OpenStudy (lindsey11x16):

cos^x(sec^2x-1) what pythagoream identiy to use?

hartnn (hartnn):

first multiply them out. like a(b-c) = ab-ac so whats cos^x(sec^2x-1) = ... - ... ?

OpenStudy (anonymous):

yeahhhhh right hartnn

OpenStudy (lindsey11x16):

cos^2x(sec^2x) - cos^2x-1

hartnn (hartnn):

yup, and you know what cos x * sec x = .. ?

hartnn (hartnn):

and -1 won't come.

hartnn (hartnn):

\(\large \cos^2x(\sec^2x-1)=\cos^2x\sec^2x-\cos^2x = ...?\)

hartnn (hartnn):

can u simplify that ^ ? use the fact that cos x = 1/ sec x.

OpenStudy (lindsey11x16):

should i use factorization in this

hartnn (hartnn):

nopes. since, cos x = 1/ sec x what is cos x * sec x = ... ?

OpenStudy (lindsey11x16):

should i turn secant into cos

hartnn (hartnn):

yes., do that...what u get ?

OpenStudy (lindsey11x16):